OpenStudy (anonymous):

Find the number of moles for each mass of substance 15.3 g berryllium

OpenStudy (trancenova):

Use this formula :) moles(n) = mass(m) / molar mass (M) You have mass (15.3g), Molar mass is found on the periodic table (9.012g/mol). (If you put the answer I will give you a medal :) )

OpenStudy (anonymous):

Oops its 53.8 g

OpenStudy (trancenova):

Lol, still just calculate with 53.8g instead of 15.3g :)

OpenStudy (anonymous):

so is it 53.8g Be* 1 mole be/18.024g Be

OpenStudy (anonymous):

so far is that okay

OpenStudy (trancenova):

Hmm no, n (moles) = 53.8/9.012 = 5.970

OpenStudy (anonymous):

Oh the formula that other similar examples were going by this formula, so confusing

OpenStudy (trancenova):

Well its the same formula you just have to re-arrange it according to the question :)

OpenStudy (anonymous):

ooh

OpenStudy (anonymous):

so moles is equal to 5.970 is that it?

OpenStudy (trancenova):

Yup :)

OpenStudy (anonymous):

So do I use the same formula with 6.79 g oxygen

OpenStudy (trancenova):

You have to use the same moles(n) = mass(m) / molar mass (M) But the molar mass will be different

OpenStudy (trancenova):

The molar mass is found under the element on the periodic table :) http://www.ptable.com/

OpenStudy (anonymous):

So i calculated it comes out to be 0.212mole 02

OpenStudy (trancenova):

Yepp! :)